OpenStudy (anonymous):

what is the simplified form of 4y^2-1/2y^2+3y-2

OpenStudy (anonymous):

4y^2-1/2y^2+3y-2 \[\frac{(2y+1)(2y-1)}{2y^2+4y-y-2}\] \[\frac{(2y+1)(2y-1)}{(2y-1)(y+2)}\] \[\frac{(2y+1)}{(y+2)}\]

OpenStudy (anonymous):

medal plx